OpenStudy (anonymous):

PLEASE HELP !!!!! Which set of line segments with the given lengths form a right triangle? 9, 27, 81 10, 24, 38 14, 16, 22 15, 36, 39

OpenStudy (anonymous):

15.36,39

OpenStudy (anonymous):

(15x15)+(36x36)=(39x39) 225+1296=1521 1521=1521 therefore D forms a right triangle

OpenStudy (ybarrap):

If the square root of the sum square of any two equals the square of the last, then these are segments of a right triangle.That is \( \sqrt{a^2 +b^2}=c^2\), then these are segments of a right triangle.

OpenStudy (anonymous):

thanks so much guys

OpenStudy (anonymous):

welcome

OpenStudy (ybarrap):

np